OpenStudy (anonymous):

7x - 3y = 4 2x - 4y = 1 Which of the following system of equations is not equal to the system of equations shown above? 28x - 12y = 16 and -6x + 12y = -3 14x - 6y = 4 and -14x + 28y = 1 -28x + 12y = -16 and 28x - 56y = 14

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 7x - 3y = 4 }\) \(\color{#000000 }{ \displaystyle 2x - 4y = 1 }\) I would multiply the first equation times 4, and second equation times -3.

OpenStudy (anonymous):

How did you come up with that?

OpenStudy (solomonzelman):

I want to make the y-coeffients equivalent, but one of them negative. I want to do this, so that the y goes away when I add the equations.

OpenStudy (solomonzelman):

is this reasonable, or are you still confused?

OpenStudy (anonymous):

Ohhh

OpenStudy (anonymous):

This is very very reasonable... these are new equation..

OpenStudy (solomonzelman):

ok, when system of equations do you get when you multiply as I directed?

OpenStudy (anonymous):

28x -12y = 16 -6x + 12y = -3

OpenStudy (solomonzelman):

\(\color{#000000 }{ \displaystyle 28x - 12y = 16 }\) \(\color{#000000 }{ \displaystyle -6x +12 y =-3 }\) good!

OpenStudy (anonymous):

Oh look! That's the first option!

OpenStudy (solomonzelman):

now, add the equations:)

OpenStudy (anonymous):

I think I understand it now :D

OpenStudy (solomonzelman):

ok, good:)

OpenStudy (anonymous):

So...

OpenStudy (anonymous):

now let's refer to the question again

OpenStudy (anonymous):

7x - 3y = 4 2x - 4y = 1 Which of the following system of equations is not equal to the system of equations shown above? 28x - 12y = 16 and -6x + 12y = -3 14x - 6y = 4 and -14x + 28y = 1 -28x + 12y = -16 and 28x - 56y = 14

OpenStudy (anonymous):

It's saying, which is NOT equal

OpenStudy (anonymous):

Is it the middle one?

OpenStudy (solomonzelman):

well, if you were to attempt to eliminate the x, what would you do with your system?

OpenStudy (solomonzelman):

times what number would you multiply equation 1 and equation 2?

OpenStudy (anonymous):

I would.... multiply the first equation by 2 and the second by ..

OpenStudy (anonymous):

I'm not sure :(

OpenStudy (anonymous):

and the second by 7?

OpenStudy (solomonzelman):

yes, but, by -7, so that you have 14x in equation 1 and -14x in equation 2 (and addition will cancel them)

OpenStudy (anonymous):

OHHH! lol you're a good teacher. What grade are u in?

OpenStudy (solomonzelman):

I am in college

OpenStudy (anonymous):

That makes sense. I am in grade 10 lol

OpenStudy (anonymous):

Thank you !! :D

OpenStudy (solomonzelman):

You are welcome!